OpenStudy (anonymous):

please help more questions in reply area 5. Lijuan is solving the equation x2 – x = 14 with the quadratic formula. What values should she use for a, b, and c? (Points : 4) a = –1, b = 1, c = –14 a = 1, b = –1, c = –14 a = –1, b = 1, c = 14 a = 1, b = –1, c = 14

OpenStudy (anonymous):

What number would you have to add to both sides to complete the square? x2 + 12x = –6 (Points : 4) 144 36 –36 –144 Question 3. 3. Solve by using the perfect squares method. x2 + 12x + 36 = 0 (Points : 4) –18 –6 6 18 this is it please help :D

OpenStudy (anonymous):

@LaceyPokemon17 @Anonymous_Hacker

OpenStudy (perl):

\[\large ax^2+bx+c = 0 \\ \large {\color{Red} 1}x^2-{\color{Red} 1}x-{\color{Red}{14}} =0\]

OpenStudy (anonymous):

i dont understand that stuff .-. tbh i cant join classes on my computer and it doesnt explain anything i just need the answers because i hvae about 9 more test in algbra :}